# Gravitational odd-parity perturbation of a Horndeski hairy black hole: quasinormal mode and parameter constraint

### v2

@@ III.2 Numerical results
In order to explicitly see the effect of the Horndeski hair on the QNMs, we plot the frequencies with first three nodes as functions of hairy charge for $\ell=2,3,4$ modes in FIG.2. For all the modes except $\ell=n=2$ (the blue solid curves), the imaginary part increases while the real part decreases as the scaled hairy parameter becomes smaller, but the imaginary part, related with the damping rates by $-Im(\omega)\sim 1/\tau$, is always negative. For one thing, this makes the odd perturbation always decay and therefore the black hole can remain stable in the range of $\tilde{Q}\in\{-1,1\}$. For another, this causes the odd perturbation to have a longer damping time. This is consistent with the phenomena we can observe in time domain (FIG.3), that the ringing stage of the perturbation with $\tilde{Q}=-1$ is always more lasting than that for other parameters. In addition, for the tendency of the real part $Re(\omega)$, it can be obviously seen that the decreasing of $\tilde{Q}$ can suppress the oscillation of the odd perturbation, which corresponds to sparse ringing for smaller $\tilde{Q}$ in time domain. Regarding the effects of angular momentum and overtone, we find that the influence of $\ell$ on the $Im(\omega)$ is much slighter than that of $n$. Moreover, it is interesting to notice that as $\tilde{Q}$ becomes negative enough, the long-lived modes with the same node are replaced by the $\ell>2$ modes, instead of the typical $\ell=2$ mode in the usual scenario. For the higher overtone, the $Im(\omega)$ of $\ell=2$ mode is highly suppressed by the negative hairy charge. In particular, both $Im(\omega)$ and $Re(\omega)$ of the node $n=2$ for $\ell=2$ mode will slightly change the tendency at $\tilde{Q}\sim-0.8$, indicating novel effects of the hairy charge on the higher overtones.
@@
Figure 2: The dimensionless QNMs as the functions of the hairy parameter $Q$ for various modes. The different curves are denoted by the angular momenta $\ell=$ 2 (solid),3 (dashed),4 (dotted) and nodes $n=$ 0 (green),1 (orange),2 (blue), respectively.
@@
Figure 4: The fundamental QNMs as the functions of the hairy parameter $Q$, for the perturbed massless scalar ($s=0,\ell=0$), Dirac ($s=1/2,\ell=1$), electromagnetic ($s=1,\ell=1$) and GW odd ($s=2,\ell=2$) fields.
@@
In our previous work Yang et al. (2024), we have addressed the QNMs of massless scalar field, electromagnetic field and Dirac field around the Horndeski hairy black hole, so we collect those fundamental QNMs frequencies and the current QNMs of gravitational odd-parity perturbation into FIG.4 and shall discuss the effect of hairy parameter on how the QNMs depend on the field’s spin. Our results show that for positive $\tilde{Q}$, the GW odd perturbation has the longest damping time and the strongest oscillation in their fundamental ($n=0$) and lowest-lying modes. However, as $\tilde{Q}$ is approaching $-1$, the GW odd perturbation turns to the shortest-lived field. We argue that this phenomenon may attribute to the superluminal propagation of the odd perturbation especially for more negative hairy charge (see FIG. 1).
@@
From the investigation on the QNMs, we find that the effect of the hairy parameter apparently differs the BH in Horndeski gravity from GR by significantly influencing the lifetime and the oscillation of gravitational odd perturbation. This will help to recognize and distinguish the hairy black hole in constructing the ringdown wave pattern and even constrain the hairy parameter in the ringdown observation. An interesting question comes follow and says that, base on the yielding QNMs data, to what extent can we constrain the horndeski BH hair according to the ringdown signal captured by GW interferometers. We will partly discuss this issue in next section.
IV Hairy parameter estimation by detection of ringdown wave
In this section, we will determine the measurement errors of the hairy parameter $\tilde{Q}$ by the GW detectors, which can be estimated by computing the Fisher information matrix (FIM) Berti et al. (2006a). To this end, we shall follow the settings in Lahoz and Noller (2023) and start with a popular template of ringdown signals in GR
@@ IV Hairy parameter estimation by detection of ringdown wave
$\omega_{QNM}=2\pi f_{\ell mn}+\frac{i}{\tau_{\ell mn}},\quad\mathscr{Q}_{\ell mn}=\pi f_{\ell mn}\tau_{\ell mn}$ (30)
where $f_{\ell mn}$ are the oscillation frequency, and $\tau_{\ell mn}$ are related with the damping time. To begin with the error estimation under the detection of above waveform (27), we need to define a useful inner product(|)($\cdot$|$\cdot$)( ⋅ | ⋅ ), weighted by a noise spectral density $S_{h}(v)$ of a detector, between the signals $h_{1}(t)$ and $h_{2}(t)$:
@@
with $\tilde{h}(v)$ defined by the Fourier transformation of the waveform $h(t)$ Therefore the signal-to-noise-ratio (SNR) can now be defined as the self-inner product of a single waveform
@@
For a preliminary work to perform parameter estimation, one usually impose some simplifying assumptions Berti et al. (2006a):
@@
$A^{\times}_{\ell mn}$ $=$ $A^{+}_{\ell mn}N_{\times}$ (35)
$\phi^{\times}_{\ell mn}$ $=$ $\phi^{+}_{\ell mn}+\phi^{0}_{\ell mn}$ (36)
where the constant phase $\phi^{0}_{\ell mn}$ and the numerical factor $N_{\times}$ can be fixed because they are addressed to be less dependent on error Berti et al. (2006a).
We shall then concentrate our interest on the measurement error of the hairy parameter $\tilde{Q}$ encoded within the QNMs frequencies $\omega_{QNM}$ such that we have $f_{\ell mn}(\tilde{Q})$ and $\mathscr{Q}_{\ell mn}(\tilde{Q})$. Since we are focusing on constraining a single parameter, so the Fisher matrix reduces to the “Fisher scalar” $\Gamma_{\tilde{Q}}=(\partial_{\tilde{Q}}h|\partial_{\tilde{Q}}h)$. Recalling the aforementioned assumptions, the derivative on the waveform, $\partial_{\tilde{Q}}h$, is reflected in the derivatives on $\partial_{\tilde{Q}}f_{\ell mn}$ and $\partial_{\tilde{Q}}\mathscr{Q}_{\ell mn}$. Since our QNMs frequencies related with $\{f_{\ell mn},\mathscr{Q}_{\ell mn}\}$ are obtained discretely, we shall do these derivatives numerically in computing the Fisher scalar as well as the errors of the hairy parameter. Consequently, for a fixed hairy parameter $\tilde{Q}_{0}$, the rescaled error $\sigma_{\tilde{Q_{0}}}$ defined in this way describes the deviation from its best estimated value by the detectors. We shall evaluate the rescaled error of the hairy parameter by considering the signals composing of a single mode and two modes, respectively.
@@ IV.1 Single-mode parameter estimation
Since our black hole is static and spherically symmetric, the odd gravitational perturbation equation here is independent of the azmuthal $m$ Ganguly et al. (2018), so we obviate $m$ and consider the detecting signals are contributed only by a single mode with
@@
Here we set $A^{+}=\frac{A^{2}}{1+N_{\times}^{2}}$, and $\{N_{\times}=1,\phi^{+}=\phi^{0}=0\}$ for simplicity, and apply the formulas (3134) laid out above, then the error (re-scaled by SNR) for a single-mode signal reads
@@
where the prime represents the derivative with respect to $\tilde{Q}$. As it was addressed in Vallisneri (2008) that $\Gamma^{-1}$ corresponds to the Cramer-Rao bound, which is a lower bound for the variance error of any unbiased parameter. So in our case the error $\rho\sigma_{n\ell}(\tilde{Q})$ possesses the similar capability that describes a lower bound variance (re-scaled by SNR) on determining the source parameter $\tilde{Q}$. We compute the tendency of error as a function of hairy parameter $\tilde{Q}$ for different single mode $h_{n\ell}$. As shown in FIG. 5, each curve describes a scaled lowest variance, which would be the best extent where the detector can constrain the hairy parameter in our setup. It can be clearly seen that the error bound of the hairy parameter becomes deeper as $\ell$ increases but shallower as $n$ increases, while the effect from the hairy parameter is relatively slight. In the figure, we also plot the errors for a large-$\mathscr{Q}$ approximation $\rho^{2}\sigma^{2}_{n\ell}\simeq\frac{f_{n\ell}^{2}}{2f_{n\ell}^{\prime 2}\mathscr{Q}_{n\ell}^{2}}$ which is directly derived from (39). A direct comparison reveals that this approximation works better for the error estimation of the fundamental mode.
Figure 5: The SNR-scaled error for single-mode $h_{n\ell}$ measurements of hairy parameter $\tilde{Q}$, with various modes. The dotted curves are the errors from a large-$\mathscr{Q}$ approximation.
@@ IV.2 Two-modes parameter estimation
The waveform mixed two different modes could be more significant in fitting astrophysical gravitational wave sources and testing the general no-hair theorem Berti et al. (2006a, b); Barack and Cutler (2004), so in this part we would like to discuss how the constraint response in considering the mixed waveforms. To model such signal, we consider the superposition of two QNMs with different $\ell$ and $n$, which can be written as
@@
where for simplicity we have labeled the mode $(\ell,n)$ as “$a$” and $(\ell^{\prime},n^{\prime})$ as “$b$”. In dealing with the product of the angular scalar between two different modes, we follow the assumption from Berti et al. (2006a) that two close enough QNMs directly give
@@
The appearing of delta symbol in (43) and (44) makes the hairy parameter estimation divide into two cases:
@@
Here $\sigma_{nab}$ denotes the variance of the mixed waveforms which have a same overtone but different angular numbers, for example, $\sigma_{023}$ means the mix of the waveforms $h_{02}$ and $h_{03}$. In this case, similar to that in single mode estimation, the mixed mode with same overtone but higher angular numbers possess lower errors as shown in the left panel of FIG.6. In order to discuss the subtle connections between the single mode and mixed mode estimation, we plot the middle and right panels of FIG.6 from which we can read off the following features: (1) when $h_{b}=h_{n\ell^{\prime}}$ dominates the mixed mode, i.e., $A_{a}/A_{b}<1$, the constraint becomes stronger and quickly approaches the single mode $h_{b}$ itself (see the green curves). (2) when $h_{a}=h_{n\ell}$ dominates the mixed mode, i.e., $A_{a}/A_{b}>1$, the constraint becomes weaker. Therefore, our results show that the measurement error from the mixed modes is rather sensitive to the their relative amplitude.
Figure 6: The SNR-scaled error for two-modes $\{h_{a},h_{b}\}=\{h_{n\ell},h_{n\ell^{\prime}}\}$ measurements of hairy parameter $\tilde{Q}$, the single modes $h_{a}$ and $h_{b}$ share a same overtone but different angular numbers. In the left plot, we have $(n,\ell,\ell^{\prime})=(1,2,3)$ (blue), $(1,3,4)$ (orange), $(0,2,3)$ (green) and $(0,3,4)$ (red), respectively. The middle and right plots present the error for the mixed modes $(n,\ell,\ell^{\prime})=(1,2,3)$ (middle) and $(1,3,4)$ (right), but with different amplitude ratio $A_{a}/A_{b}$ between the two modes. The dotted curves denotes the single-mode errors.
@@
where $\sigma_{\ell ab}$ denotes the variance of the mixed mode with the same angular number but different overtones, for example, $\sigma_{201}$ mixes the waveforms $h_{02}$ and $h_{12}$. The Key results are depicted in the left panel of FIG.7, showing that the mixed modes with same angular momentum but smaller overtones exhibit lower error. To more explicitly see the mixing effects, we plot the middle and right panels. The middle and right panels indicate that when the fundamental mode $h_{a}=h_{n\ell}$ dominates the mixing, the overall error becomes smaller (see the green curves in middle and right panels of FIG.7 ), but if the next-to-leading overtone mode $h_{b}=h_{n^{\prime}\ell}$ dominates the mixed modes, the error becomes larger. In addition, due to the existence of the interacting term $\Gamma^{(c)}_{\tilde{Q}}$, we also observe intriguing phenomena that the detection potential from the mixed modes is not always better than that from the single mode, and it is significantly affected by the amplitude ratio $A_{a}/A_{b}$ of the mixing mode as well as the hairy parameter.
Figure 7: The SNR-scaled error for two-modes $\{h_{a},h_{b}\}=\{h_{n\ell},h_{n^{\prime}\ell}\}$ measurements of hairy parameter $\tilde{Q}$, the single modes $h_{a}$ and $h_{b}$ share a same angular number but different overtones. In the left plot, we have $(\ell,n,n^{\prime})=(2,1,2)$ (blue), $(3,1,2)$ (orange), $(2,0,1)$ (green), $(3,0,1)$ (red), respectively. In the middle and right plots, we select the error with $(\ell,n,n^{\prime})=(2,0,1)$ (middle) and $(2,1,2)$ (right) and check the effect of the amplitude ratio $A_{a}/A_{b}=1/5$ of the two modes. The dotted curves are the single-mode errors.

### v3

@@ III.2 Numerical results
In order to explicitly see the effect of the Horndeski hair on the QNMs, we plot the frequencies with the first three nodes as functions of hairy charge for $\ell=2,3,4$ modes in FIG.2. For all the modes except $\ell=n=2$ (the blue solid curves), the imaginary part increases while the real part decreases as the scaled hair parameter becomes smaller, but the imaginary part, related to the damping rates by $-Im(\omega)\sim 1/\tau$, is always negative. For one thing, this causes the odd perturbation to always decay, and therefore the black hole can remain stable in the range of $\tilde{Q}\in\{-1,1\}$. For another, this causes the odd perturbation to have a longer damping time. This is consistent with the phenomena we can observe in time domain (FIG.3), that the ringing stage of the perturbation with $\tilde{Q}=-1$ is always lasting longer than that for other parameters. In addition, for the tendency of the real part $Re(\omega)$, it can be obviously seen that the decreasing of $\tilde{Q}$ can suppress the oscillation of the odd perturbation, which corresponds to a sparse oscillations for smaller $\tilde{Q}$ in the time domain. Regarding the effects of angular momentum and overtone, we find that the influence of $\ell$ on $Im(\omega)$ is much weaker than that of $n$. Moreover, it is interesting to notice that as $\tilde{Q}$ becomes sufficiently negative, the long-lived modes with the same node are replaced by the $\ell>2$ modes, instead of the typical $\ell=2$ mode in the usual GR scenario. For higher overtones, the $Im(\omega)$ of $\ell=2$ mode is highly suppressed by the negative hairy charge. In particular, both $Im(\omega)$ and $Re(\omega)$ for the mode $\ell=n=2$ will slightly change their trend at $\tilde{Q}\sim-0.8$, indicating novel effects of the hairy charge on the higher overtones.
@@
Figure 2: The dimensionless QNMs as functions of the hair parameter $\tilde{Q}$ for various modes. The different curves are denoted by the angular momenta $\ell=$ 2 (solid),3 (dashed),4 (dotted) and nodes $n=$ 0 (green),1 (orange),2 (blue), respectively.
@@
Figure 4: The fundamental QNMs as functions of the hair parameter $\tilde{Q}$, for the perturbed massless scalar ($s=0,\ell=0$), Dirac ($s=1/2,\ell=1$), electromagnetic ($s=1,\ell=1$) and GW odd ($s=2,\ell=2$) fields.
@@
In our previous work Yang:2023lcm , we addressed the QNMs of massless scalar field, electromagnetic field and Dirac field around the Horndeski hairy black hole, so we collect those fundamental QNMs frequencies and the current QNMs of odd perturbation into FIG.4, and shall discuss the effect of hair parameter on how the QNMs depend on the field’s spin. Our results show that for positive $\tilde{Q}$, the odd perturbation has the longest damping time and the strongest oscillation in their fundamental ($n=0$) and lowest-lying modes. However, as $\tilde{Q}$ is approaching $-1$, the odd perturbation turns into the shortest-lived field. To further understand the abnormal behavior of the odd perturbation at the analytic level, we shall analyze its dynamics in the eikonal ($\ell>>1$) limit, in which the potential of the standard wave equation (34) reduces to
$V_{e}(r)=\ell^{2}\frac{f}{r^{2}},$ (39)
and the QNM frequency can be analytically extracted as PhysRevD.35.3621
$\omega_{\text{eikonal}}=\sqrt{V_{e}(r_{0})}-i(n+\frac{1}{2})\sqrt{\frac{-1}{2V_{e}(r_{0})}\left(\frac{\mathrm{d^{2}}}{\mathrm{d}\tilde{r}^{2}_{*}}V_{e}\right)_{r=r_{0}}}\,,$ (40)
where $r_{0}$ is the position where $V_{e}(r)$ reaches its maximum  [^1] This process of extracting QNMs frequencies in eikonal limit is not proper for the master equation (33) because it is not the standard wave equation for WKB analysis.. It is noted that the effective potential (39) is the same as those for various massless matter field perturbations Yang:2023lcm for which in the eikonal limit, their QNM frequencies can also be calculated by the geometric-optics approximation formula Cardoso:2008bp
$\omega_{\text{GOA}}=\ell~{}\Omega_{c}-i(n+\frac{1}{2})|\lambda_{LE}|,$ (41)
$\text{with}\quad\Omega_{c}=\sqrt{\frac{f}{r_{c}}}\quad\text{and}\quad|\lambda_{LE}|=\sqrt{-\frac{r^{2}_{c}}{2f_{c}}\left(\frac{\mathrm{d^{2}}}{\mathrm{d}\bar{r}^{2}_{*}}\frac{f}{r^{2}}\right)_{r=r_{c}}}.$ (42)
Here $r_{c}$ is the radius of the null circular orbit, given by the positive root of the equation $02f(r)-rf^{\prime}(r)=$ 0, so we have $r_{c}=r_{0}$. Carefully comparing (40) and (41), we find that their real part are consistent, but the imaginary parts mismatch because the tortoise coordinates, $\tilde{r}_{*}$ and $\bar{r}_{*}$, are different. This implies that though the link between the null geodesics and quasinormal modes for the massless extrenal field perturbations is fulfilled, it is violated for the gravitational odd perturbation in the current theory. Note that a similar violation was also found in Einstein-Lovelock theory, in which the effective potential of the gravitational perturbation in the eikonal limit was different from radial potential of the null particle Konoplya:2017wot . Additionally, we plot FIG. 5 to show how the Horndeski hair affect the mismatch between $\omega_{\text{eikonal}}$ and $\omega_{\text{GOA}}$ for the odd perturbation.
Figure 5: The fundamental QNMs as functions of the hair parameter $\tilde{Q}$. Here we have fixed $\ell=40$. The blue curve represents the QNM obtained from the exact WKB formula (40) for the odd perturbation in the $-\ell$ limit. The orange curve represents the QNM obtained from the geometric-optics approximation formula (41), which is the QNM for scalar, Dirac and electromagnetic field perturbations degenerated to in the $-\ell$ limit. The mismatch of the imaginary part between $\omega_{\text{eikonal}}$ and $\omega_{\text{GOA}}$ obviously arises for nonvanishing Horndeski hair.
@@
From the investigation on the QNMs, we find that the effect of the hair parameter apparently differs the BH in Horndeski gravity from GR by significantly influencing the lifetime and the oscillation of gravitational odd perturbation. This will help to recognize and distinguish the hairy black hole in constructing the ringdown wave pattern and even constrain the hair parameter in the ringdown observation. An interesting question comes follow and says that, based on the yielding QNM data, to what extent can we constrain the horndeski BH hair according to the ringdown signal captured by GW interferometers. We will partly discuss this issue in next section.
IV Hair parameter estimation by detection of ringdown wave
In this section, we will determine the measurement errors of the hair parameter $\tilde{Q}$ by the GW detectors, which can be estimated by computing the Fisher information matrix (FIM) Berti:2005ys . Considering that the realistic GW events are usually related to the rotating black holes, the studies based on QNM spectrum from non rotating may not be applied in practice. However, the preliminary analysis on non-rotating scenario is still important to be carried out because it encodes the leading order information for sufficiently slow rotation cases. To this end, we start with a popular template of ringdown signals in GR
@@ IV Hair parameter estimation by detection of ringdown wave
$\omega_{\text{QNM}}=2\pi f_{\ell mn}+\frac{i}{\tau_{\ell mn}},\quad\mathscr{Q}_{\ell mn}=\pi f_{\ell mn}\tau_{\ell mn},$ (46)
where $f_{\ell mn}$ are the oscillation frequency, and $\tau_{\ell mn}$ are related to the damping time. It is noted that this template based on the ringdown QNM frequencies can really generate a complete ringdown GW waveform in GR for which the odd and even parity perturbations have the same spectrum. But in scalar-tensor theory the isospectrality was found to be broken Li:2023ulk , and other polarization modes, such as longitudinal and breathing modes, can arise due to the scalar field Eardley:1973br . Therefore, the template we utilize here is not realistic to generate a complete ringdown waveform in this specific Horndeski theory. Nevertheless, since the odd perturbation is decoupled from the scalar field, so here we assume the related ringdown GW can be treated separately and described by the aforementioned template for a preliminary study. Similar assumption was also taken in Lahoz:2023csk . To begin with the error estimation under the detection of above waveform (43), we need to define a useful inner product(|)($\cdot$|$\cdot$)( ⋅ | ⋅ ), weighted by a noise spectral density $S_{h}(v)$ of a detector, between the signals $h_{1}(t)$ and $h_{2}(t)$:
@@
with $\tilde{h}(v)$ defined by the Fourier transformation of the waveform $h(t)$ Therefore, the signal-to-noise-ratio (SNR) can now be defined as the self-inner product of a single waveform
@@
For a preliminary work to perform parameter estimation, one usually impose some simplifying assumptions Berti:2005ys :
@@
$A^{\times}_{\ell mn}$ $=$ $A^{+}_{\ell mn}N_{\times},$ (51)
$\phi^{\times}_{\ell mn}$ $=$ $\phi^{+}_{\ell mn}+\phi^{0}_{\ell mn},$ (52)
where the constant phase $\phi^{0}_{\ell mn}$ and the numerical factor $N_{\times}$ can be fixed because they are addressed to be less dependent on the error Berti:2005ys .
We shall then concentrate our interest on the measurement error of the hair parameter $\tilde{Q}$ encoded within the QNMs frequencies $\omega_{QNM}$ such that we have $f_{\ell mn}(\tilde{Q})$ and $\mathscr{Q}_{\ell mn}(\tilde{Q})$. Since we are focusing on constraining a single parameter, so the Fisher matrix reduces to a “Fisher scalar” $\Gamma_{\tilde{Q}}=(\partial_{\tilde{Q}}h|\partial_{\tilde{Q}}h)$. Recalling the aforementioned assumptions, the derivative on the waveform, $\partial_{\tilde{Q}}h$, is reflected in the derivatives on $\partial_{\tilde{Q}}f_{\ell mn}$ and $\partial_{\tilde{Q}}\mathscr{Q}_{\ell mn}$. Since our QNMs frequencies related to $\{f_{\ell mn},\mathscr{Q}_{\ell mn}\}$ are obtained discretely, we shall do these derivatives numerically in computing the Fisher scalar as well as the errors of the hair parameter. Consequently, for a fixed hair parameter $\tilde{Q}_{0}$, the rescaled error $\sigma_{\tilde{Q_{0}}}$ defined in this way describes the deviation from its best estimated value by the detectors. We shall evaluate the rescaled error of the hair parameter by considering the signals composing of a single mode and two modes, respectively.
@@ IV.1 Single-mode parameter estimation
Since our black hole is static and spherically symmetric, the odd gravitational perturbation equation here is independent of the azmuthal $m$ Ganguly:2017ort , so we obviate $m$ and consider the detecting signals are contributed only by a single mode with
@@
Here we set $A^{+}=\frac{A^{2}}{1+N_{\times}^{2}}$, and $\{N_{\times}=1,\phi^{+}=\phi^{0}=0\}$ for simplicity, and apply the formulas (4750) laid out above. Then the error (re-scaled by SNR) for a single-mode signal reads
@@
where the prime represents the derivative with respect to $\tilde{Q}$.
As it was addressed in Vallisneri:2007ev that $\Gamma^{-1}$ corresponds to the Cramer-Rao bound, which is a lower bound for the variance error of any unbiased parameter. So in our case the error $\rho\sigma_{n\ell}(\tilde{Q})$ possesses the similar capability that describes a lower bound variance (re-scaled by SNR) on determining the source parameter $\tilde{Q}$. We compute the tendency of error as a function of hair parameter $\tilde{Q}$ for different single mode $h_{n\ell}$. As shown in FIG. 6, each curve describes a scaled lowest variance, which would be the best extent where the detector can constrain the hair parameter in our setup. It can be clearly seen that the error bound of the hair parameter becomes deeper as $\ell$ increases but shallower as $n$ increases, while the effect from the hair parameter is relatively slight. In the figure, we also plot the errors for a large-$\mathscr{Q}$ approximation $\rho^{2}\sigma^{2}_{n\ell}\simeq\frac{f_{n\ell}^{2}}{2f_{n\ell}^{\prime 2}\mathscr{Q}_{n\ell}^{2}}$ which is directly derived from (55). A direct comparison reveals that this approximation works better for the error estimation of the fundamental mode.
Figure 6: The SNR-scaled error for single-mode $h_{n\ell}$ measurements of hair parameter $\tilde{Q}$, with various modes. The dotted curves are the errors from the large-$\mathscr{Q}$ approximation.
@@ IV.2 Two-modes parameter estimation
The waveform mixed two different modes could be more significant in fitting astrophysical gravitational waves sources and testing the general no-hair theorem Berti:2005ys ; Berti:2006hb ; Barack:2003fp , so in this part we would like to discuss how the constraint response in considering the mixed waveforms. To model such signal, we consider the superposition of two QNMs with different $\ell$ and $n$, which can be written as
@@
where for simplicity we have labeled the mode $(\ell,n)$ as “$a$” and $(\ell^{\prime},n^{\prime})$ as “$b$”. In dealing with the product of the angular scalar between two different modes, we follow the assumption from Berti:2005ys that two close enough QNMs directly gives
@@
The appearance of a delta symbol in (59) and (60) makes the hair parameter estimation divide into two cases:
@@
Here $\sigma_{nab}$ denotes the variance of the mixed waveforms which have a same overtone but different angular numbers, for example, $\sigma_{023}$ means the mix of the waveforms $h_{02}$ and $h_{03}$. In this case, similar to the one in the single mode estimation, the mixed mode with the same overtone but higher angular numbers exhibits lower error as shown in the left panel of FIG.7. In order to discuss the subtle connections between the single mode and mixed mode estimation, we plot the middle and right panels of FIG.7 from which we can read off the following features: (1) when $h_{b}=h_{n\ell^{\prime}}$ dominates the mixed mode, i.e., $A_{a}/A_{b}<1$, the constraint becomes stronger and quickly approaches the single mode $h_{b}$ itself (see the green curves). (2) when $h_{a}=h_{n\ell}$ dominates the mixed mode, i.e., $A_{a}/A_{b}>1$, the constraint becomes weaker. Therefore, our results show that the measurement error from the mixed modes is rather sensitive to their relative amplitude.
Figure 7: The SNR-scaled error for two-modes $\{h_{a},h_{b}\}=\{h_{n\ell},h_{n\ell^{\prime}}\}$ measurements of hair parameter $\tilde{Q}$, the single modes $h_{a}$ and $h_{b}$ share a same overtone but different angular numbers. In the left plot, we have $(n,\ell,\ell^{\prime})=(1,2,3)$ (blue), $(1,3,4)$ (orange), $(0,2,3)$ (green) and $(0,3,4)$ (red), respectively. The middle and right plots present the error for the mixed modes $(n,\ell,\ell^{\prime})=(1,2,3)$ (middle) and $(1,3,4)$ (right), but with different amplitude ratio $A_{a}/A_{b}$ between the two modes. The dotted curves denotes the single-mode errors.
Figure 8: The SNR-scaled error for two-modes $\{h_{a},h_{b}\}=\{h_{n\ell},h_{n^{\prime}\ell}\}$ measurements of hair parameter $\tilde{Q}$, the single modes $h_{a}$ and $h_{b}$ share a same angular number but different overtones. In the left plot, we have $(\ell,n,n^{\prime})=(2,1,2)$ (blue), $(3,1,2)$ (orange), $(2,0,1)$ (green), $(3,0,1)$ (red), respectively. In the middle and right plots, we select the error with $(\ell,n,n^{\prime})=(2,0,1)$ (middle) and $(2,1,2)$ (right) and check the effect of the amplitude ratio $A_{a}/A_{b}=1/5$ of the two modes. The dotted curves are the single-mode errors.
@@
where $\sigma_{\ell ab}$ denotes the variance of the mixed mode with the same angular number but different overtones, for example, $\sigma_{201}$ mixes the waveforms $h_{02}$ and $h_{12}$. The Key results are depicted in the left panel of FIG.8, showing that the mixed modes with same angular momentum but smaller overtones exhibit lower errors. To more explicitly see the mixing effects, we plot the middle and right panels. The middle and right panels indicate that when the fundamental mode $h_{a}=h_{n\ell}$ dominates the mixing, the overall error becomes smaller (see the green curves in middle and right panels of FIG.8 ), but if the next-to-leading overtone mode $h_{b}=h_{n^{\prime}\ell}$ dominates the mixed modes, the error becomes larger. In addition, due to the existence of the interacting term $\Gamma^{(c)}_{\tilde{Q}}$, we also observe intriguing phenomena that the detection potential from the mixed modes is not always better than that from the single mode, and it is significantly affected by the amplitude ratio $A_{a}/A_{b}$ of the mixing mode as well as the hair parameter.